OpenStudy (anonymous):

consider the equation: 2cos(4x) +1 = 0 find all degree solutions Help please

OpenStudy (anonymous):

2cos(4x)=-1 cos4x=-1/2 4x=c0s^-1(-1/2)

OpenStudy (anonymous):

x=cos^-1(-1/2)/4

OpenStudy (dumbcow):

^^ what he said now what is arccos of -1/2 ? use unit circle

OpenStudy (anonymous):

cos inverse

OpenStudy (anonymous):

I get confused when it comes to the unit circle

OpenStudy (anonymous):

\[x=(\cos ^{-1}(-1/2))/4\]

OpenStudy (anonymous):

x=-(cos1/2)/4 have u got it now

OpenStudy (anonymous):

Yes thanks :)

OpenStudy (dumbcow):

\[x = \frac{\pi}{6}, \frac{\pi}{3}\]

OpenStudy (anonymous):

can you explain how you got that